OpenStudy (elcna):

Help please? I'll fan and medal :) What is (-m)^-3n if m=2 and n=-24? *3 *-3 *4 *-4 Can someone explain how to do this to me because I keep getting -193/8 and I don't know if that's completely wrong or not?

jimthompson5910 (jim_thompson5910):

do you mean \[\Large (-m)^{-3n}\] ??

OpenStudy (elcna):

No, the 'n' variable isn't part of the exponent @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so this? \[\Large (-m)^{-3}n\]

OpenStudy (elcna):

Yep

jimthompson5910 (jim_thompson5910):

We replace m with 2 and n with -24 and simplify \[\Large (-m)^{-3}n\] \[\Large (-2)^{-3}(-24)\] \[\Large \frac{1}{(-2)^3}*(-24)\] \[\Large \frac{1}{-8}*(-24)\] \[\Large \frac{-24}{-8}\] \[\Large 3\] Therefore, \[\Large (-m)^{-3}n=3\] when m = 2 and n = -24

OpenStudy (anonymous):

You got it right elca

jimthompson5910 (jim_thompson5910):

On step 3, I used the rule \[\Large x^{-k} = \frac{1}{x^k}\]

OpenStudy (elcna):

Thank you!! :) @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome